HUBBLE SPACE TELESCOPE DAILY REPORT #1165 PERIOD COVERED: 0700 EST Daily Status Report as of 173/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxies 3C309.1 and 3C410.0 as part of a continuum survey and to study the galaxies' morphology. On the first set of observations, two images were received and displayed. Both were reported to look okay. The second set of observations was executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5419 (A Search for Low Luminosity Active Galactic Nuclei in Nearby Galaxies) The WF/PC-2 was used to make observations of galaxy NGC-3992, which is known to have a ultraviolet bright, low-luminosity, active galactic nuclei. Three images were received and displayed. They were reported to look okay. 1.3 Completed FOS/BL 5362 (A Time Resolved UV Study of the Wind Outflow) The Faint Object Spectrograph (Blue Side) was used to make observations of the cataclysmic binary star V795-Herculis which is known to make spectrographic changes in the order of a few minutes and other longer periods. Ten observations were received and displayed. They were reported to look okay, with the maximum counts in groups 3, 10, and 8 of the 1X3, 6x2, and 3X3 peakups, respectfully. 1.4 Completed WF/PC-2 5369 (HST Medium Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. Four observations were received and displayed, and they were reported to look okay. More observations still remain on Science Tape Recorder #1, and will be analyzed later. No anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1: Flats Part 2) The WF/PC-2 was used to make visual flat calibration observations. Six images were received and displayed. Five were reported to be nominal, and one was underexposed. More observations still remain on Science Tape Recorder #1, and will be analyzed later. No anomalies were noted. 1.6 Completed FOC/96 5330 (High Resolution Mapping of Pluto's Albedo Distribution) The Faint Object Camera was used to make observations of the planet Pluto and its moon Charon in two bandpasses. The observations were executed as scheduled, and no anomalies were noted. The observations still remain on Science Tape Recorder #1, and will be analyzed later. 1.7 Completed WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earthflats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed FOS/BL 5539 (Cycle 4 FOS Photometric Monitor) The Faint Object Spectrograph (Blue Side) used standard star BD+28D4211 to calibrate the absolute sensitivity of its detectors. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled :10 Successful :10 2.3 Significant HSTARs: None 2.4 Operations Notes: There were no significant problems in any HST subsystem during the last twenty-four hours. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: There will be a HST/RFSOC/Second TDRS Ground Terminal (STGT) Test conducted tomorrow (Thursday, June 23) in preparation for the upcoming Spacecraft Test scheduled for Aug. 6, 1994. /DMH -END-